# AttributionScanner: A Visual Analytics System for Model Validation with Metadata-Free Slice Finding

### v3

@@ IV-B Bird Category Classifier Validation - Detecting Bias
$\bullet$ Does the model exhibit bias? To answer this key question, experts start by investigating the underperformed slices. From the Slice Table (Fig. 7  A), they sort slices by ascending order of accuracy and select the worst-performed slice_38. The coordinated information provided by Attribution Mosaic and model attribution heatmaps highlights a spurious correlation problem—the model uses water backgrounds to classify birds (Fig. 10).
@@
The experts annotate this slice as “spurious”, and AttributionScanner automatically propagates this annotation. They verify the propagation correctness on neighborhood slices and annotate slice_41 as “spurious” with “water backgrounds” (Fig. 10). Moreover, the experts identify an underperformed slice_32 that is not clustered with the current ones and is given a high Spuriousness possibility. They investigate it and verify that the model utilizes “Spurious feature: land backgrounds”, to predict bird classes. Through the Attribution Mosaic in the confusion matrix form (Fig. 7  D), they find such spurious correlations result in many false negatives (Fig. 7  D1), where the model uses land backgrounds to mistakenly predict many “waterbirds” as “landbirds”.
$\bullet$ Is the detected bias prevalent across all slices? Why or why not? ML experts are interested in determining whether the detected bias is pervasive throughout the dataset. By analyzing slices that are distant from the annotated ones in the Attribution Mosaic and are assigned with low Spuriousness possibilities, they discovered that the farthest neighbors, namely slice_34 and slice_5, correspond to core features. This suggests that the model can correctly capture bird regions in these slices (Fig. 7  B  C), which raises a follow-up “why” question. To understand in what circumstances when the model fails. They browse the original images from slice_32 (spurious feature) and slice_34 (core feature), respectively, as shown in Fig. 7  D and  C. They find slice_32 has very similar land backgrounds and very different birds, while on the other hand, the birds’ appearance of slice_34 is very consistent. This finding explains why this biased model can successfully capture the core features from slice_34 but fails at slice_32—greater similarities in the representation space indicate stronger features. In other words, core features in slice_34 are strong enough to support the model’s robustness against bias. Such insights are helpful in improving model robustness and have been further studied by ML experts [5].
@@ V Evaluation
In this section, we conduct both quantitative and qualitative evaluations to validate whether AttributionScanner can indeed leverage human insights to enhance vision models’ performance while reducing their reliance on spurious features.
@@ V-C Experts Feedback
Other comments.    Our experts provided feedback on additional aspects. One expert suggested that we directly display predictions and labels via text in each confusion matrix panel. “It would be better if you show ‘label: waterbirds; prediction: landbirds’ rather than ‘TN’”. Another expert discussed the potential of applying our system to validate multi-classification models. Following our brief explanation of how AttributionScanner produces results based on the model’s predicted class without restrictions on the number of classes, the expert commented, “Your system seems to be fully adaptable. And I would love to see this implementation in future work.”
@@ Acknowledgments
This research is supported in part by the National Science Foundation via grant No. IIS-2427770, and by the University of California (UC) Multicampus Research Programs and Initiatives (MRPI) grant and the UC Climate Action Initiative grant.
[…]
@@
  • Sec. A-V provides the discussion and future work.
@@ A-IV. Label Spreading Algorithm
A-V. Discussion and Future Work

### v4

@@ IV-B Bird Category Classifier Validation - Detecting Bias
$\bullet$ Does the model exhibit bias? To answer this key question, experts start by investigating the underperformed slices. From the Slice Table (Fig. 7  A), they sort slices by ascending order of accuracy and select the worst-performed slice_38. The coordinated information provided by Attribution Mosaic and model attribution heatmaps highlights a spurious correlation problem—the model uses water backgrounds to classify birds (Fig. 10). The experts annotate this slice as “spurious”, and AttributionScanner automatically propagates this annotation. They verify the propagation correctness on neighborhood slices and annotate slice_41 as “spurious” with “water backgrounds” (Fig. 10). Moreover, the experts identify an underperformed slice_32 that is not clustered with the current ones and is given a high Spuriousness possibility. They investigate it and verify that the model utilizes “Spurious feature: land backgrounds”, to predict bird classes. Through the Attribution Mosaic in the confusion matrix form (Fig. 7  D), they find such spurious correlations result in many false negatives (Fig. 7  D1), where the model uses land backgrounds to mistakenly predict many “waterbirds” as “landbirds”.
$\bullet$ Is the detected bias prevalent across all slices? Why or why not? ML experts are interested in determining whether the detected bias is pervasive throughout the dataset.
@@
By analyzing slices that are distant from the annotated ones in the Attribution Mosaic and are assigned with low Spuriousness possibilities, they discovered that the farthest neighbors, namely slice_34 and slice_5, correspond to core features. This suggests that the model can correctly capture bird regions in these slices (Fig. 7  B  C), which raises a follow-up “why” question. To understand in what circumstances when the model fails. They browse the original images from slice_32 (spurious feature) and slice_34 (core feature), respectively, as shown in Fig. 7  D and  C. They find slice_32 has very similar land backgrounds and very different birds, while on the other hand, the birds’ appearance of slice_34 is very consistent. This finding explains why this biased model can successfully capture the core features from slice_34 but fails at slice_32—greater similarities in the representation space indicate stronger features. In other words, core features in slice_34 are strong enough to support the model’s robustness against bias. Such insights are helpful in improving model robustness and have been further studied by ML experts [5].
@@ Acknowledgments
This research is supported in part by the National Science Foundation via grant No. IIS-2427770, the National Institute of Health under grant no. P41 EB032840, and the University of California (UC) Multicampus Research Programs and Initiatives (MRPI) grant and the UC Climate Action Initiative grant.
@@
  • Sec. A-V provides additional feedback from our experts.
  • Sec. A-VI provides the discussion and future work.
@@ A-IV. Label Spreading Algorithm
A-V. Additional experts feedback
Our experts provided feedback on additional aspects. One expert suggested that we directly display predictions and labels via text in each confusion matrix panel. “It would be better if you show ‘label: waterbirds; prediction: landbirds’ rather than ‘TN’”. Another expert discussed the potential of applying our system to validate multi-classification models. Following our brief explanation of how AttributionScanner produces results based on the model’s predicted class without restrictions on the number of classes, the expert commented, “Your system seems to be fully adaptable. And I would love to see this implementation in future work.”
A-VI. Discussion and Future Work
@@
Extending to object detection and segmentation tasks.     Although our system is designed primarily for image classification, its core methodology—data slice generation and visualization techniques—can be extended to tasks such as object detection and segmentation. These extensions would require adjustments in slice definition and model performance metrics, but the attribution-based approach remains applicable. For instance, object detection models often provide region-specific attribution scores, which can be used to generate localized data slices. Similarly, segmentation tasks could leverage attribution maps to group pixels into coherent slices for validation. Future work will explore such extensions to broaden the applicability of our framework to a wider range of vision tasks and multimodal datasets.
Scalability and computational efficiency improvements.     Our current system is optimized for handling large-scale datasets efficiently. Feature inversion, despite being an optimization-based method, is applied at the slice level rather than to individual instances, allowing for parallelized processing. Empirical results on CelebA (with $202k$ images) show that our feature inversion generates 50 slice visualizations in under 3 minutes using an Nvidia RTX 3090 GPU. Additionally, components such as UMAP, K-Means, and label propagation are designed to handle high-dimensional data at scale.